OpenStudy (anonymous):

2. The distance traveled by a falling object is given by the formula d = 0.5gt2 where d = distance, g = the force of gravity, and t = time. Solve this equation for g, and use your formula to determine the force of gravity if a baseball takes 10 seconds to hit the ground after being dropped from a height of 480 feet. Show all steps in your work.

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

This one is just like the other one, I figured out the equation but I don't know how to plug in the numbers.

OpenStudy (anonymous):

The formula is: 2d/t2 = g

OpenStudy (anonymous):

Would the height = distance?

OpenStudy (anonymous):

Nvm I got the answer of g = 9.6

OpenStudy (anonymous):

Can anyone check that?

OpenStudy (jack1):

d = 0.5gt2 where d = distance = 480 feet g = the force of gravity, t = time = 10 seconds \[ d = \frac 12 gt^2 \] \[ 480 = \frac 12 g \times 10^2 \] \[ 480 = \frac 12 g \times 100 \] \[ 480 = g \times \frac 12 \times 100 \] \[ 480 = g \times 50 \] \[ \frac{480}{50} = g \]

OpenStudy (anonymous):

Where's your equation to solve for "g"?

OpenStudy (anonymous):

Whatever, your answer is the same as mine.

OpenStudy (jack1):

yep, urs is perfect

OpenStudy (anonymous):

Can you help me check another?

OpenStudy (jack1):

sure dude, shoot

OpenStudy (anonymous):

Two boys want to use a seesaw and they need to move the seesaw so that their weights will balance out. The formula is given by w1 • d1 = w2 • d2 where w1 = weight of the first boy, d1 = distance of the first boy from the fulcrum, w2 = weight of the second boy, and d2 = distance of the second boy from the fulcrum. Rewrite the formula to solve for d2. Show all steps in your work.

OpenStudy (anonymous):

My answer was w1 x d1/w2 = d2

OpenStudy (jack1):

yep, correct

OpenStudy (anonymous):

Thanks.

OpenStudy (jack1):

welcomes dude

OpenStudy (anonymous):

I need help with another I'll close this and reopen another.